Alfalfa data (a2010) taken during sep12
showed long recovery times with alfa beam1b when the radars were
- oct09: Mitsubishi amplifiers outfitted with leds to solve
radar long recovery times
- apr12: alfa amplifiers replaced with nrao amplifiers. The
amps have led's installed
- sep12: alfalfa runs and we see long recovery times on
beam1b when the radars are present.
- 10oct12: checked led currents. 3milliamps (as they were
supposed to be).
Data setup and processing:
- a2010 does 600 second drift scans at the meridian using
- 100 Mhz bandwidth centered around 1380 Mhz.is taken.
- Data is sampled at 1 sec intervals.
- The total power for each one second is computed over the
100 mhz. The computation tries to not include frequency
channels with rfi.
The plots show the total power vs
time for alfalfa (.ps)
- 600 seconds of data from 1 drift scan. The sample rate is
once per second.
- Each color is a different beam.
- The vertical display units are Tsys. A vertical offset has
been added to each beam for display.
- Top frame is polA, bottom frame is polB.
- Page 1: 600 seconds of data from 1 drift scan.
- The positive going spikes are the 12 second radars.
- beam 1b (bottom frame, red trace) has negative going
- This band has gone into compression
- You can see that it takes many seconds for the value
- Page 2: a blowup showing seconds 180 to 380.
- You can see the long recovery time for beam1b
- beam 0b looks a like it may have a little delay.
- long recovery times were seen on 10sep12 with the a2010
- 10oct12: denis checked the led currents. They were 3
milliamps (as the notes say they should be).
- Previously the long recovery times came because the
Mitsubishi amps didn't have leds.
- We don't know whether the current problem is in the
amps, or somewhere else.
a2130 cal deflection data has shown pol1b cal
deflection jumping around a lot more than the other beams.
- This occurred during:
- nov11-dec11.. this is before the new amps were
- So maybe this problem is separate from the slow
- or maybe the slow recovery in 1b is not happening in
- After the new amps were installed in apr12, i never
checked for the slow recovery time of the new amps on the
telescope with the radars present. Looking back at the a2048
data may-sep12 shows:
- baseline variations in bm1b: 31may12,
- long recovery from radars: 02jun,25aug,
- So it looks like bm1b has been acting up for awhile.